# Studies of stationary features in jets: 3C 279 quasar I. On-sky scattering and dynamics

### v1

@@ 7 Summary
We investigate the dynamics of the quasi-stationary component (QSC) in the flat-spectrum radio quasar 3C 279 using 15 GHz VLBA monitoring data from the MOJAVE program. The QSC is identified in 161 epochs spanning over 27 years, with its median position located at a separation of $\sim 0.35$ mas from the radio core (or projected distance of $\approx 2.2$ pc).
Studies of the on-sky distribution of QSC positions have proven to be a valuable tool for analysing changes in the jet direction and cone opening angle. Our analysis reveals that the jet axis has changed direction by approximately $20\degr$ over this 27-year period. During this time, the jet axis has gradually shifted southward, showing signs of nutation. The apparent cone angle of the QSC in good agreement with the apparent jet cone angle inferred from the jet width and the motion of components downstream of the QSC and it varies over time, with a mean value of about $16\degr\pm 2\degr$, corresponding to a deprojected intrinsic cone angle of $0.3\degr\pm 0.03\degr$, assuming a jet viewing angle of $1\degr$. The scatter of QSC positions across the jet axis reflects the maximum amplitude of the transverse waves propagating through the QSC.
The on-sky distribution of the QSC’s apparent displacement vectors reveals an excess in both number and magnitude along the jet axis, suggesting strong anisotropic displacements of the core in the jet direction. To quantify this effect, we applied a method that disentangles apparent displacements and allows for statistical estimation of the intrinsic apparent displacements of both the QSC and the core. We found that the intrinsic rmsc of core displacements exceeds the intrinsic rmss of the QSC by a factor of 1.9. Using the estimated mean intrinsic displacement of the QSC and the standard deviation of its displacements, we estimated the mean speed of the QSC as $\overline{\beta}_{\rm s}\approx 16\pm 2$. Within the seagull-on-wave model, the observed superluminal speed of the QSC provides evidence for relativistic transverse waves propagating through it. We further estimated the approximate mean speed of the transverse wave ($\Gamma_{\rm w}\approx 60$) and its mean inclination angle ($\overline{\varphi}\approx 0.3\degr$) relative to the jet axis.
The characteristics of the QSC’s reversing trajectory and its classification enable the determination of the amplitudes and frequencies of these relativistic transverse waves. By analyzing the smoothed and refined trajectory of the QSC, we estimated a wave frequency of $\sim 0.5$ yr-1 and a mean wave amplitude of $\overline{A}_{\rm w}\approx 0.06\pm 0.005$ mas ($\approx 0.38$ pc), which in fact represent the lower and upper limits, respectively. The estimate of $\overline{A}_{\rm w}$ is consistent with the mean maximum wave amplitude derived from the transverse scatter of QSC positions, $\overline{A}_{\rm w}<\overline{A}_{\rm w,max}=0.1$ mas.
@@
We conclude that studying the dynamics of the QSC is a valuable tool for probing the presence of relativistic transverse waves in the jets of BL Lac objects and FSRQs, as well as for determining their key characteristics, including frequency, speed, amplitude, and magnetic energy.
@@ Acknowledgements.
We thank the MOJAVE team for comments on the manuscript. The VLBA is a facility of the National Radio Astronomy Observatory, a facility of the National Science Foundation that is operated under cooperative agreement with Associated Universities, Inc. This research has made use of data from the MOJAVE database that is maintained by the MOJAVE team (Lister et al. 2018). The research was supported by the YSU, in the frames of the internal grant.

### v2

@@ 7 Summary
We have investigated the dynamics of the QSC in the FSRQ 3C 279 using 15 GHz VLBA monitoring data from the MOJAVE programme. The QSC is identified in 161 epochs spanning over 27 years, with its median position located at a separation of $\sim 0.35$ mas from the radio core (or projected distance of $\approx 2.2$ pc).
Studies of the on-sky distribution of QSC positions have proven to be a valuable tool for analysing changes in the jet direction and cone opening angle. Our analysis reveals that the jet axis has changed direction by approximately $20\degr$ over this 27-year period. During this time, the jet axis has gradually shifted southwards, showing signs of nutation. The apparent cone angle of the QSC is in good agreement with the apparent jet cone angle inferred from the jet width and the motion of components downstream of the QSC and it varies over time, with a mean value of about $16\degr\pm 2\degr$, corresponding to a deprojected intrinsic cone angle of $0.3\degr\pm 0.03\degr$, assuming a jet viewing angle of $1\degr$. The scatter of QSC positions across the jet axis reflects the maximum amplitude of the transverse waves propagating through the QSC.
The on-sky distribution of the QSC’s apparent displacement vectors reveals an excess in both number and magnitude along the jet axis, suggesting strong anisotropic displacements of the core in the jet direction. To quantify this effect, we applied a method that disentangles apparent displacements and allows for statistical estimation of the intrinsic apparent displacements of both the QSC and the core. We found that the intrinsic rmsc of core displacements exceeds the intrinsic rmss of the QSC by a factor of $\approx 1.6$. Using the estimated mean intrinsic displacement of the QSC and the standard deviation of its displacements, we estimated the mean speed of the QSC as $\overline{\beta}_{\rm s}\approx 10\pm 1.3$. Within the seagull-on-wave model, the observed superluminal speed of the QSC provides evidence of relativistic transverse waves propagating through it. We further estimated the approximate mean speed of the transverse wave ($\Gamma_{\rm w}\approx 60$) and its mean inclination angle ($\overline{\varphi}\approx 0.2\degr$) relative to the jet axis.
The characteristics of the QSC’s reversing trajectory and its classification enable the amplitudes and frequencies of these relativistic transverse waves to be determined. By analysing the smoothed and refined trajectory of the QSC, we estimated a wave frequency of $\sim 0.5$ yr-1 and a mean wave amplitude of $\overline{A}_{\rm w}\approx 0.06\pm 0.005$ mas ($\approx 0.38$ pc), which in fact represent the lower and upper limits, respectively. The estimate of $\overline{A}_{\rm w}$ is consistent with the mean maximum wave amplitude derived from the transverse scatter of QSC positions, $\overline{A}_{\rm w}<\overline{A}_{\rm w,max}=0.1$ mas.
@@
We conclude that studying the dynamics of the QSC is a valuable tool for probing the presence of relativistic transverse waves in the jets of BL Lac objects and FSRQs. It also enables the determination of their key characteristics, including frequency, speed, amplitude, and magnetic energy.
@@ Acknowledgements.
We thank the MOJAVE team for comments on the manuscript and the anonymous referee for careful reading and valuable suggestions. The VLBA is a facility of the National Radio Astronomy Observatory, a facility of the National Science Foundation that is operated under cooperative agreement with Associated Universities, Inc. This research has made use of data from the MOJAVE database that is maintained by the MOJAVE team (lister18). The research was supported by the YSU, in the frames of the internal grant.